# A conjecture on descents, inversions and the weak order on Coxeter groups

### v2

@@ 4.1. Type $B$ permutations, inversions and standardization
For the proof that Conjecture 1 holds in type $B$, we need the interpretation of inversions and descents on injective words over the alphabet $[\![\bar{n},n]\!]^{*}$ as discussed in §3.1 for the alphabet $[n]$. More precisely, any element $\sigma\in W_{n}$ is an element of the symmetric group $\mathcal{S}_{[\![\bar{n},n]\!]^{*}}$ and, therefore, has a set of inversion $\operatorname{invs}(\sigma)$ as an injective words on the alphabet $[\![\bar{n},n]\!]^{*}$ (see Eq.(1)):
$\operatorname{invs}(\sigma)=\{(\sigma_{i},\sigma_{j})\mid i,j\in[\![\bar{n},n]\!]^{*},\ i<j,\ \sigma_{i}>\sigma_{j}\}.$
@@
  • $(\sigma_{\bar{i}},\sigma_{i})\in\operatorname{invs}(\sigma)\iff(\sigma_{\bar{i}}\ \sigma_{i})\in T(\sigma)$;
  • $((\sigma_{\bar{j}}\ \sigma_{\bar{i}})\in\operatorname{invs}(\sigma)\iff(\sigma_{i}\ \sigma_{j})\in\operatorname{invs}(\sigma)\iff(\sigma_{\bar{i}}\ \sigma_{\bar{j}})(\sigma_{i}\ \sigma_{j})\in T(\sigma)$.
@@
(6) $T(\sigma)=\{(a\ |a|)\mid(a,|a|)\in\operatorname{invs}(\sigma)\}\sqcup\{(\bar{a}\ \bar{b})(a\ b)\mid(a\ b)\in\operatorname{invs}(\sigma)\}.$
For simplicity, we say that $(a,b)\in\operatorname{invs}(\sigma)$ is an inversion of $\sigma$. Of course, as transpositions, $(\bar{a}\ \bar{b})=(\bar{b}\ \bar{a})$, but $(\bar{a},\bar{b})\not=(\bar{b},\bar{a}))$ with the notations used for $\operatorname{invs}(\sigma)$. The right descent sets is:
@@
The inversions of $\operatorname{invs}(\sigma)$ are:
$\operatorname{invs}(\sigma)$ $T(\sigma)$ $\operatorname{invs}(\sigma)$ $T(\sigma)$
$(\bar{1},1)$ (1¯ $(\bar{1}\ 1)$ $(\bar{4},1)$, $(\bar{1},4)$ (4¯ 1)(1¯ $(\bar{4}\ 1)(\bar{1}\ 4)$
$(\bar{3},3)$ (3¯ $(\bar{3}\ 3)$ $(\bar{4},3)$, $(\bar{3},4)$ (4¯ 3)(3¯ $(\bar{4}\ 3)(\bar{3}\ 4)$
$(\bar{4},4)$ (4¯ $(\bar{4}\ 4)$ $(\bar{4},\bar{2})$, $(2,4)$ (4¯2¯)(2 $(\bar{4}\ \bar{2})(2\ 4)$
$(\bar{4},5)$, $(\bar{5},4)$ (4¯ 5)(5¯ $(\bar{4}\ 5)(\bar{5}\ 4)$ $(5,6)$, $(\bar{6},\bar{5})$ (5 $(5\ 6)(\bar{6}\ \bar{5})$
$(\bar{4},2)$, $(\bar{2},4)$ (4¯ 2)(2¯ $(\bar{4}\ 2)(\bar{2}\ 4)$ $(2,6)$, $(\bar{6},\bar{2})$ (2 $(2\ 6)(\bar{6}\ \bar{2})$
$(\bar{4},6)$, $(\bar{6},4)$ (4¯ 6)(6¯ $(\bar{4}\ 6)(\bar{6}\ 4)$ $(2,3)$, $(\bar{3},\bar{2})$ (2 $(2\ 3)(\bar{3}\ \bar{2})$
$(\bar{4},\bar{3})$, $(3,4)$ (4¯3¯)(3 $(\bar{4}\ \bar{3})(3\ 4)$ $(\bar{3},\bar{1})$, $(1,3)$ (3¯1¯)(1 $(\bar{3}\ \bar{1})(1\ 3)$
$(\bar{4},\bar{1})$ , $(1,4)$ (4¯1¯)(1 $(\bar{4}\ \bar{1})(1\ 4)$ $(\bar{3},1)$, $(\bar{1},3)$ (3¯ 1)(1¯ $(\bar{3}\ 1)(\bar{1}\ 3)$
The set $\operatorname{invs}(\sigma)$ has cardinality $29$ whereas $\ell(\sigma)=16=|T(\sigma)|$.
Even though the sets $\operatorname{invs}(\sigma)$ and $T(\sigma)$ are different, we obtain directly from Eq.(6) the following characterization of bipartitions of elements of $W_{n}$.
@@
Let $w$ be an element of $W_{n}$, then $\{u,v\}$ forms a bipartition of $\sigma$ if and only if $\operatorname{invs}(w)=\operatorname{invs}(u)\sqcup\operatorname{invs}(v)$.
@@
Let $I$ be a set of tuples $(a,b)$ with $a,b\in[\![\bar{n},n]\!]^{*}$ for some $n\geq 1$ and $a<b$. Then $I=\operatorname{invs}(\sigma)$ of a certain word $\sigma$ corresponding to an element of $W_{n}$ if and only if it satisfies the following for all $a,b,c\in[\![\bar{n},n]\!]^{*}$ with $a<b<c$
@@
In particular, Conditions (1) and (2) are the same as in type $A$, only Condition (3) is added to take into account the symmetry of the signed permutation. In other words, an element of $W_{n}$ is only a permutation of size $2n$ (on the set $[\![\bar{n},n]\!]^{*}$) with some extra symmetry conditions.
@@
Let $w$ be the word corresponding to an element of $I_{n}$. Let $I$ be a subset of $[\![\bar{n},n]\!]^{*}$. We write $\bar{I}$ the set $\{\bar{a};a\in I\}$. We say that $I$ is symmetric if $I=\bar{I}$. On the contrary, we say that $I$ is antisymmetric if $I\cap\bar{I}=\varnothing$. Now
  • for $I$ an antisymmetric subset of $[\![\bar{n},n]\!]^{*}$ [^3] This definition is actually valid on any subset, but we only need in this article it for antisymmetric subsets., then $\operatorname{std}(w_{|I})$ is the classical standardization of the word $w$ restricted to $I$. The result is a permutation (of type $A$) of size $|I|$.
  • for $I$ a symmetric subset of $[\![\bar{n},n]\!]^{*}$, we write $\operatorname{std}^{B}(w_{|I})$ the type $B$ standardization of the word $w$ restricted to $I$, i.e., this is the classical standardization except that we use the alphabet $[\![\bar{k},k]\!]^{*}$ where $k=\frac{|I|}{2}$. The result is then an element of type $W_{k}$.
@@
For the remainding of this section, we identify the elements of $B_{n}$ with their words as signed permutations. The proof of Theorem 1.5 in the case of type $B$ is very similar to type $A$. The difficulty lies in finding a recursive decomposition of type $B$ permutations that satisfies the same properties as in type $A$.
@@
Let $w$ be an element of $W_{n}$ with $n\geq 2$. The decreased word of $w$ is obtained by removing both the letters $n$ and $\bar{n}$ from $w$ such that $\tilde{w}$ is now a word corresponding to an element of $W_{n-1}$.
We use the same notation as in type $A$ as it is always clear from the context which decreasing is applied. For example, if $w=4\bar{5}\bar{2}\bar{6}31|\bar{1}\bar{3}625\bar{4}$, then $\tilde{w}=4\bar{5}\bar{2}31|\bar{1}\bar{3}25\bar{4}$. We state the following result, which proof is similar to the one of type $A$ and is left to the reader.
@@
Let $\{u,v\}$ be a bipartition of an element $w\in W_{n}$ with $n>1$, then $\{\tilde{u},\tilde{v}\}$ is a bipartition of $\tilde{w}$.
@@
$n\textrm{ is in {\em negative position} in }w\iff n\textrm{ is in p{\em ositiveposition} in }ww_{\circ}.$
Let $\{u,v\}$ be a bipartition of an element $w\in W_{n}$ with $n>1$. Assume that $n$ is in negative position in $w$. Then $(n,\bar{n})\in\operatorname{invs}(w)$. Since the roles of $u$ and $v$ are symmetric, we have without loss of generality that $n$ is in negative position in $u$ and is in positive position in $v$. Therefore $n$ is in positive position in $ww_{\circ}$, $uw_{\circ}$ and $v$. Moreover, by Proposition 2.6, $\{ww_{\circ},v\}$ is a bipartition of $uw_{\circ}$ and Conjecture 1 holds for the triplet $u,v,w$ if and only if it holds for $ww_{\circ},v,uw_{\circ}$. The following proposition resume these facts.
@@
We define the right letters of $w$, $\mathcal{R}(w)$, the forgotten letters of $w$, $\mathcal{F}(w)$, and the left letters of $w$, $\mathcal{L}(w)$ as follows:
$\mathcal{R}(w)$ :={a($:=\{a\mid(a,n)\in\operatorname{invs}(w)\}\cup\{n\};$
$\mathcal{F}(w)$ $:=\overline{\mathcal{R}(w)};$
$\mathcal{L}(w)$ :=$:=\{a\mid a\not\in\mathcal{R}(w)\cup\mathcal{F}(w)\}.$
The left (resp. right) words of $w$, $u$, and $v$ are respectively:
$w_{\mathcal{L}}$ $=\operatorname{std}^{B}(w_{|\mathcal{L}(w)}),$ $w_{\mathcal{R}}$ $=\operatorname{std}(w_{|\mathcal{R}(w)})$
$u_{\mathcal{L}}$ $=\operatorname{std}^{B}(u_{|\mathcal{L}(w)}),$ $u_{\mathcal{R}}$ $=\operatorname{std}(u_{|\mathcal{R}(w)})$
$v_{\mathcal{L}}$ $=\operatorname{std}^{B}(v_{|\mathcal{L}(w)}),$ $v_{\mathcal{R}}$ $=\operatorname{std}(v_{|\mathcal{R}(w)}).$
@@
Assume that $n$ is in positive position in $w\in W$. Observe that all letters of $\mathcal{R}(w)$ are in positive positions: so $\mathcal{R}(w)$ is an antisymmetric subset of $[\![\bar{n},n]\!]^{*}$ (see Definition 4.6). On the other hand, $\mathcal{L}(w)$ is defined as the complement of a symmetric set of letters, so it is itself a symmetric subset of $[\![\bar{n},n]\!]^{*}$.
In the examples below, whenever two letters belong to the same set $\mathcal{L}$, $\mathcal{R}$, or $\mathcal{F}$, we say that they are of the same color. In the following, we often color left letters in red, right letters in blue and forgotten letters in black.
@@
We obtain ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{R}(w)}=\{\bar{4},2,5,6\}$, $\mathcal{F}(w)=\{4,\bar{2},\bar{5},\bar{6}\}$, and ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathcal{L}(w)}=\{\bar{3},\bar{1},1,3\}$. Therefore:
@@
$w_{\mathcal{L}}$ $=\operatorname{std}^{B}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}31|\bar{1}\bar{3}})=21|\bar{1}\bar{2},$ $w_{\mathcal{R}}$ $=\operatorname{std}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}625\bar{4}})=4231,$
$u_{\mathcal{L}}$ $=\operatorname{std}^{B}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}3\bar{1}|1\bar{3}})=2\bar{1}|1\bar{2},$ $u_{\mathcal{R}}$ $=\operatorname{std}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}625\bar{4}})=4231,$
$v_{\mathcal{L}}$ $=\operatorname{std}^{B}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bar{3}1|\bar{1}3})=\bar{2}1|\bar{1}2,$ $v_{\mathcal{R}}$ $=\operatorname{std}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bar{4}256})=1234.$
@@
Consider $w=1\bar{4}\bar{3}\bar{6}\bar{2}\bar{5}|52634\bar{1}$. Observe that $n=6$ is in positive position in $w$. The elements $u=\bar{4}\bar{3}\bar{6}\bar{2}\bar{5}\bar{1}|152634$ and $v=1\bar{6}\bar{5}\bar{4}\bar{3}\bar{2}|23456\bar{1}$ form a bipartition of $w$. We obtain ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{R}(w)}=\{\bar{1},3,4,6\}$, $\mathcal{F}(w)=\{1,\bar{3},\bar{4},\bar{6}\}$, and ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathcal{L}(w)}=\{\bar{5},\bar{2},2,5\}$. Hence:
@@
$w_{\mathcal{L}}$ $=\operatorname{std}^{B}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bar{2}\bar{5}|52})=\bar{1}\bar{2}|21,$ $w_{\mathcal{R}}$ $=\operatorname{std}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}634\bar{1}})=4231,$
$u_{\mathcal{L}}$ $=\operatorname{std}^{B}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bar{2}\bar{5}|52})=\bar{1}\bar{2}|21,$ $u_{\mathcal{R}}$ $=\operatorname{std}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bar{1}634})=1423,$
$v_{\mathcal{L}}$ $=\operatorname{std}^{B}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bar{5}\bar{2}|25})=\bar{2}\bar{1}|12,$ $v_{\mathcal{R}}$ $=\operatorname{std}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}346\bar{1}})=2341.$
@@
Consider $w=2\bar{1}\bar{3}\bar{4}5\bar{6}|6\bar{5}431\bar{2}$,. Observe that $n=6$ is in positive position in $w$. The elements $u=2\bar{1}\bar{4}\bar{3}5\bar{6}|6\bar{5}341\bar{2}$ and $v=\bar{6}\bar{5}\bar{3}\bar{4}\bar{2}\bar{1}|124356$ form a bipartition of $w$. As $n$ is in position $1$, ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{R}(w)}$ contains all letters in positive positions while $\mathcal{F}(w)$ contains all letters in negative positions and ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathcal{L}(w)}$ is empty. We get
@@
$w_{\mathcal{R}}$ $=\operatorname{std}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6\bar{5}431\bar{2}})=615432,$
$u_{\mathcal{R}}$ $=\operatorname{std}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6\bar{5}341\bar{2}})=614532,$
$v_{\mathcal{R}}$ $=\operatorname{std}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bar{5}\bar{2}1436})=123546.$

### v3

@@ 4.1. Type $B$-permutations, inversions and standardization
For the proof that Conjecture 1 holds in type $B$, we need the interpretation of inversions and descents on injective words over the alphabet $\bar{n},n]\!]^{*}$ as discussed in §3.1 for the alphabet $[n]$. More precisely, any element $\sigma\in W_{n}$ is an element of the symmetric group $\mathcal{S}_{[\![\bar{n},n]\!]^{*}}$ and, therefore, has a set of inversion $\operatorname{Inv}(\sigma)$ as an injective words on the alphabet $\bar{n},n]\!]^{*}$ (see Eq.(1)):
$\operatorname{Inv}(\sigma)=\{(\sigma_{i},\sigma_{j})\mid i,j\in[\![\bar{n},n]\!]^{*},\ i<j,\ \sigma_{i}>\sigma_{j}\}.$
@@
  • $(\sigma_{\bar{i}},\sigma_{i})\in\operatorname{Inv}(\sigma)\iff(\sigma_{\bar{i}}\ \sigma_{i})\in T(\sigma)$;
  • $(\sigma_{\bar{j}}\ \sigma_{\bar{i}})\in\operatorname{Inv}(\sigma)\iff(\sigma_{i}\ \sigma_{j})\in\operatorname{Inv}(\sigma)\iff(\sigma_{\bar{i}}\ \sigma_{\bar{j}})(\sigma_{i}\ \sigma_{j})\in T(\sigma)$.
@@
(6) $T(\sigma)=\{(a\ |a|)\mid(a,|a|)\in\operatorname{Inv}(\sigma)\}\sqcup\{(\bar{a}\ \bar{b})(a\ b)\mid(a\ b)\in\operatorname{Inv}(\sigma)\}.$
For simplicity, we say that $(a,b)\in\operatorname{Inv}(\sigma)$ is an inversion of $\sigma$. Of course, as transpositions, $(\bar{a}\ \bar{b})=(\bar{b}\ \bar{a})$, but $(\bar{a},\bar{b})\not=(\bar{b},\bar{a}))$ with the notations used for $\operatorname{Inv}(\sigma)$. The right descent sets is:
@@
The inversions of $\operatorname{Inv}(\sigma)$ are:
$\operatorname{Inv}(\sigma)$ $T(\sigma)$ $\operatorname{Inv}(\sigma)$ $T(\sigma)$
$(\bar{1},1)$ (1¯ 1)($\bar${1}\ 1) $(\bar{4},1)$, $(\bar{1},4)$ (4¯ 1)(1¯ 4)($\bar${4}\ 1)($\bar${1}\ 4)
$(\bar{3},3)$ (3¯ 3)($\bar${3}\ 3) $(\bar{4},3)$, $(\bar{3},4)$ (4¯ 3)(3¯ 4)($\bar${4}\ 3)($\bar${3}\ 4)
$(\bar{4},4)$ (4¯ 4)($\bar${4}\ 4) $(\bar{4},\bar{2})$, $(2,4)$ (4¯2¯)(2 4)($\bar${4}\ $\bar${2})(2\ 4)
$(\bar{4},5)$, $(\bar{5},4)$ (4¯ 5)(5¯ 4)($\bar${4}\ 5)($\bar${5}\ 4) $(5,6)$, $(\bar{6},\bar{5})$ (5 6)(6¯5¯)(5\ 6)($\bar${6}\ $\bar${5})
$(\bar{4},2)$, $(\bar{2},4)$ (4¯ 2)(2¯ 4)($\bar${4}\ 2)($\bar${2}\ 4) $(2,6)$, $(\bar{6},\bar{2})$ (2 6)(6¯2¯)(2\ 6)($\bar${6}\ $\bar${2})
$(\bar{4},6)$, $(\bar{6},4)$ (4¯ 6)(6¯ 4)($\bar${4}\ 6)($\bar${6}\ 4) $(2,3)$, $(\bar{3},\bar{2})$ (2 3)(3¯2¯)(2\ 3)($\bar${3}\ $\bar${2})
$(\bar{4},\bar{3})$, $(3,4)$ (4¯3¯)(3 4)($\bar${4}\ $\bar${3})(3\ 4) $(\bar{3},\bar{1})$, $(1,3)$ (3¯1¯)(1 3)($\bar${3}\ $\bar${1})(1\ 3)
$(\bar{4},\bar{1})$ , $(1,4)$ (4¯1¯)(1 4)($\bar${4}\ $\bar${1})(1\ 4) $(\bar{3},1)$, $(\bar{1},3)$ (3¯ 1)(1¯ 3)($\bar${3}\ 1)($\bar${1}\ 3)
The set $\operatorname{Inv}(\sigma)$ has cardinality $29$ whereas $\ell(\sigma)=16=|T(\sigma)|$.
Even though the sets $\operatorname{Inv}(\sigma)$ and $T(\sigma)$ are different, we obtain directly from Eq.(6) the following characterization of bipartitions of elements of $W_{n}$.
@@
Let $w$ be an element of $W_{n}$, then $\{u,v\}$ forms a bipartition of $\sigma$ if and only if $\operatorname{Inv}(w)=\operatorname{Inv}(u)\sqcup\operatorname{Inv}(v)$.
@@
Let $I$ be a set of tuples $(a,b)$ with $a,b\in[\![\bar{n},n]\!]^{*}$ for some $n\geq 1$ and $a<b$. Then $I=\operatorname{Inv}(\sigma)$ of a certain word $\sigma$ corresponding to an element of $W_{n}$ if and only if it satisfies the following for all $a,b,c\in[\![\bar{n},n]\!]^{*}$ with $a<b<c$
@@
In particular, Conditions (1) and (2) are the same as in type $A$, only Condition (3) is added to take into account the symmetry of the signed permutation. In other words, an element of $W_{n}$ is only a permutation of size $2n$ (on the set $\bar{n},n]\!]^{*}$) with some extra symmetry conditions.
@@
Let $w$ be the word corresponding to an element of $W_{n}$. Let $I$ be a subset of $\bar{n},n]\!]^{*}$. We write $\bar{I}$ the set $\{\bar{a};a\in I\}$. We say that $I$ is symmetric if $I=\bar{I}$. On the contrary, we say that $I$ is antisymmetric if $I\cap\bar{I}=\varnothing$. Now
  • for $I$ an antisymmetric subset of $\bar{n},n]\!]^{*}$ [^4] This definition is actually valid on any subset, but we only need in this article it for antisymmetric subsets., then $\operatorname{std}(w_{|I})$ is the classical standardization of the word $w$ restricted to $I$. The result is a permutation (of type $A$) of size $|I|$.
  • for $I$ a symmetric subset of $\bar{n},n]\!]^{*}$, we write $\operatorname{std}^{B}(w_{|I})$ the type $B$ standardization of the word $w$ restricted to $I$, i.e., this is the classical standardization except that we use the alphabet $\bar{k},k]\!]^{*}$ where $k=\frac{|I|}{2}$. The result is then an element of type $W_{k}$.
@@
For the remainding of this section, we identify the elements of $W_{n}$ with their words as signed permutations. The proof of Theorem 3.1 in the case of type $B$ is very similar to type $A$. The difficulty lies in finding a recursive decomposition of type $B$ permutations that satisfies the same properties as in type $A$.
@@
Let $w$ be an element of $W_{n}$ with $n\geq 2$. The decreased word of $w$ is obtained by removing both the letters $n$ and $\bar{n}$ from $w$ such that $\widetilde{w}$ is now a word corresponding to an element of $W_{n-1}$.
We use the same notation as in type $A$ as it is always clear from the context which decreasing is applied. For example, if $w=4\bar{5}\bar{2}\bar{6}31|\bar{1}\bar{3}625\bar{4}$, then $\widetilde{w}=4\bar{5}\bar{2}31|\bar{1}\bar{3}25\bar{4}$. We state the following result, which proof is similar to the one of type $A$ and is left to the reader.
@@
Let $\{u,v\}$ be a bipartition of an element $w\in W_{n}$ with $n>1$, then $\{\widetilde{u},\widetilde{v}\}$ is a bipartition of $\widetilde{w}$.
@@
Let $\{u,v\}$ be a bipartition of an element $w\in W_{n}$ with $n>1$. Assume that $n$ is in negative position in $w$. Then $(n,\bar{n})\in\operatorname{Inv}(w)$. Since the roles of $u$ and $v$ are symmetric, we have without loss of generality that $n$ is in negative position in $u$ and is in positive position in $v$. Therefore $n$ is in positive position in $ww_{\circ}$, $uw_{\circ}$ and $v$. Moreover, by Proposition 2.6, $\{ww_{\circ},v\}$ is a bipartition of $uw_{\circ}$ and Conjecture 1 holds for the triplet $u,v,w$ if and only if it holds for $ww_{\circ},v,uw_{\circ}$. The following proposition resume these facts.
@@
We define the right letters of $w$, ${\overrightarrow{n}}(w)$, the forgotten letters of $w$, $\mathcal{F}(w)$, and the left letters of $w$, ${\overleftarrow{n}}(w)$ as follows:
${\overrightarrow{n}}(w)$ $:=\{a\mid(a,n)\in\operatorname{Inv}(w)\}\cup\{n\};$
$\mathcal{F}(w)$ $:=\overline{{\overrightarrow{n}}(w)};$
${\overleftarrow{n}}(w)$ $:=\{a\mid a\not\in{\overrightarrow{n}}(w)\cup\mathcal{F}(w)\}.$
These letters are elements of $\bar{n},n]^{*}$. The left (resp. right) words of $w$, $u$, and $v$ are respectively:
$w_{\overleftarrow{n}}$ $=\operatorname{std}^{B}(w_{|{\overleftarrow{n}}(w)}),$ $w_{\overrightarrow{n}}$ $=\operatorname{std}(w_{|{\overrightarrow{n}}(w)})$
$u_{\overleftarrow{n}}$ $=\operatorname{std}^{B}(u_{|{\overleftarrow{n}}(w)}),$ $u_{\overrightarrow{n}}$ $=\operatorname{std}(u_{|{\overrightarrow{n}}(w)})$
$v_{\overleftarrow{n}}$ $=\operatorname{std}^{B}(v_{|{\overleftarrow{n}}(w)}),$ $v_{\overrightarrow{n}}$ $=\operatorname{std}(v_{|{\overrightarrow{n}}(w)}).$
@@
Assume that $n$ is in positive position in $w\in W$. Observe that all letters of ${\overrightarrow{n}}(w)$ are in positive positions: so ${\overrightarrow{n}}(w)$ is an antisymmetric subset of $\bar{n},n]\!]^{*}$ (see Definition 4.6). On the other hand, ${\overleftarrow{n}}(w)$ is defined as the complement of a symmetric set of letters, so it is itself a symmetric subset of $\bar{n},n]\!]^{*}$.
In the examples below, whenever two letters belong to the same set $\overleftarrow{n}$}, $\overrightarrow{n}$}, or $\mathcal{F}$, we say that they are of the same color. In the following, we often color left letters in red, right letters in blue and forgotten letters in black.
@@
We obtain ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\overrightarrow{n}}(w)}=\{\bar{4},2,5,6\}$, $\mathcal{F}(w)=\{4,\bar{2},\bar{5},\bar{6}\}$, and ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{\overleftarrow{n}}(w)}=\{\bar{3},\bar{1},1,3\}$. Therefore:
@@
$w_{\overleftarrow{n}}$ $=\operatorname{std}^{B}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}31|\bar{1}\bar{3}})=21|\bar{1}\bar{2},$ $w_{\overrightarrow{n}}$ $=\operatorname{std}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}625\bar{4}})=4231,$
$u_{\overleftarrow{n}}$ $=\operatorname{std}^{B}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}3\bar{1}|1\bar{3}})=2\bar{1}|1\bar{2},$ $u_{\overrightarrow{n}}$ $=\operatorname{std}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}625\bar{4}})=4231,$
$v_{\overleftarrow{n}}$ $=\operatorname{std}^{B}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bar{3}1|\bar{1}3})=\bar{2}1|\bar{1}2,$ $v_{\overrightarrow{n}}$ $=\operatorname{std}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bar{4}256})=1234.$
@@
Consider $w=1\bar{4}\bar{3}\bar{6}\bar{2}\bar{5}|52634\bar{1}$. Observe that $n=6$ is in positive position in $w$. The elements $u=\bar{4}\bar{3}\bar{6}\bar{2}\bar{5}\bar{1}|152634$ and $v=1\bar{6}\bar{5}\bar{4}\bar{3}\bar{2}|23456\bar{1}$ form a bipartition of $w$. We obtain ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\overrightarrow{n}}(w)}=\{\bar{1},3,4,6\}$, $\mathcal{F}(w)=\{1,\bar{3},\bar{4},\bar{6}\}$, and ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{\overleftarrow{n}}(w)}=\{\bar{5},\bar{2},2,5\}$. Hence:
@@
$w_{\overleftarrow{n}}$ $=\operatorname{std}^{B}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bar{2}\bar{5}|52})=\bar{1}\bar{2}|21,$ $w_{\overrightarrow{n}}$ $=\operatorname{std}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}634\bar{1}})=4231,$
$u_{\overleftarrow{n}}$ $=\operatorname{std}^{B}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bar{2}\bar{5}|52})=\bar{1}\bar{2}|21,$ $u_{\overrightarrow{n}}$ $=\operatorname{std}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bar{1}634})=1423,$
$v_{\overleftarrow{n}}$ $=\operatorname{std}^{B}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bar{5}\bar{2}|25})=\bar{2}\bar{1}|12,$ $v_{\overrightarrow{n}}$ $=\operatorname{std}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}346\bar{1}})=2341.$
@@
Consider $w=2\bar{1}\bar{3}\bar{4}5\bar{6}|6\bar{5}431\bar{2}$,. Observe that $n=6$ is in positive position in $w$. The elements $u=2\bar{1}\bar{4}\bar{3}5\bar{6}|6\bar{5}341\bar{2}$ and $v=\bar{6}\bar{5}\bar{3}\bar{4}\bar{2}\bar{1}|124356$ form a bipartition of $w$. As $n$ is in position $1$, $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\overrightarrow{n}}(w)$} contains all letters in positive positions while $\mathcal{F}(w)$ contains all letters in negative positions and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{\overleftarrow{n}}(w)$} is empty. We get
@@
$w_{\overrightarrow{n}}$ $=\operatorname{std}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6\bar{5}431\bar{2}})=615432,$
$u_{\overrightarrow{n}}$ $=\operatorname{std}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6\bar{5}341\bar{2}})=614532,$
$v_{\overrightarrow{n}}$ $=\operatorname{std}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bar{5}\bar{2}1436})=123546.$